OpenStudy (anonymous):

Diophantus was a child for one-sixth of his life. After one-twelth more, he became a man. After one-seventh more, he married, and five years later his son was born. His son lived half as long as his father and died four years before his father. How old was Diophantus when he died?

OpenStudy (anonymous):

Please help me!!!

OpenStudy (kropot72):

You need to solve the following equation, where the life span of Diophantus is x years: \[\frac{x}{6}+\frac{x}{12}+\frac{x}{7}+5+\frac{x}{2}+4=x\]